Hi, I seem to be stuck on a problem I can only get so far on this, I would appreciate your help please. Here is the question; Business and Finance. The Randolphs used 12 more gallons of fuel oil in October than in September and twice as much oil in November as in September. If they used 132 gal for the 3 months, how much was used during each month?
I know that evenly the 3 months used 44 gal of oil fuel, but trying to figure this out is probably so simple but I keep confusing myself on this for some reason. Any help would be appreciated.
Thank You, Barb Neely

Answer by edjones(8007)

Let x be gallons of fuel oil used in September, x+12 oct, 2x nov
x+x+12+2x=132
4x+12=132
4x+12-12=132-12
4x=120
4x/4=120/4
x=30 gal sept
30+12=42 gal oct
2*30=60 gal nov
Always check
30+42+60=132
Ed